If q1=-8.2C and q2=4.9C, what is the electric potential difference Va-Vb (in gigavolts GV)?


1
Expert's answer
2021-10-03T16:16:57-0400

Let distances:

"d_1=1" m, "d_2=0.5" m

Then:

"V_a-V_b=k(q_1/d_1-q_2/d_2)=9\cdot10^9(-8.2+9.8)=14.4" GV